Circular No. 6491

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996bk IN NGC 5308 Piero Mazza and Stefano Pesci, Milan, report their visual discovery

with a 0.40-m reflector of an apparent supernova (mag 14.5–15.0) on Oct. 12.79 UT, located ∼ 20′′ southwest of the nucleus of NGC 5308. The object is not present in various photographic atlases. Nothing unusual was noted by Pesci on Aug. 15 at this location. M. Schwartz, Cascade Mountains, OR, obtained a CCD image showing the new object on Oct. 13; no star is present at this position on an image of the galaxy taken by Schwartz on June 5. J. Mackey, Werrington, England, reports the following CCD position obtained on Oct. 13.795 UT: α = 13^{h}46^{m}57^{s}.98, δ = +60^{o}58′12′′.9 (equinox 2000.0); he measured the magnitude of SN 1996bk as V = 14.05. G. Comello, Groningen, The Netherlands, reports the following visual magnitude estimates: Oct. 13.82, 13.9; 14.80, 14.1.

P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that spectra obtained by J. Huchra and L. Macri on Oct. 15.1 UT with the 1.5-m Tillinghast telescope show that this is a type-Ia supernova about a week past maximum light. CCD images taken with the Whipple Observatory 1.2-m reflector show SN 1996bk to be 10′′.5 south and 17′′.9 west of the galaxy center, at magnitude V = 15.”

PSR 0833–45 C. Flanagan, Hartebeesthoek Radio Astronomy Observatory (HRAO),

reports that a large spin-up in PSR 0833–45 (the Vela pulsar) was observed from the HRAO, the relative increase δf/f in spin-rate being +2.15(2) × 10^{−6} on Oct. 13.394 UT.

COMET C/1996 R1 (HERGENROTHER-SPAHR) Total visual magnitude estimates: Sept. 30.77 UT, 11.6 (K. Hornoch,

Lelekovice, Czech Republic, 0.35-m reflector); Oct. 8.25, 13.0 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 13.86, 12.5 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector);

SUPERNOVAE Corrigendum. Regarding the item on IAUC 6490, B. Schmidt reports

that P. Garnavich (CfA) was also observing on the MMT with R. Kirshner.

1996 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green